OpenStudy (anonymous):

Help needed please : |x-1| < 2|x+2|

OpenStudy (anonymous):

Have you tried converting the absolutes to squares?

OpenStudy (anonymous):

yes i did @lightchaste, but do i have to squares |x+2| or the whole 2|x+2|?

OpenStudy (saifoo.khan):

not 2. coz it's not in mod.

OpenStudy (anonymous):

Just the |x+2| :)

OpenStudy (anonymous):

okay, so i've got the answers in decimals. but apparently the answers i the answer scheme is in integers

OpenStudy (saifoo.khan):

What's the next step? @lightchaste

OpenStudy (anonymous):

What's the answer?

OpenStudy (anonymous):

the answer is x<-5 and x>-1

OpenStudy (saifoo.khan):

Correct.

OpenStudy (aravindg):

ys thats what twolf also says http://www.wolframalpha.com/input/?i=%7Cx-1%7C+%3C+2%7Cx%2B2%7C

OpenStudy (anonymous):

okay so @saifoo.khan, how did it get to be -5 and -1 bcs i seem to be getting decimals while doing the quadratic equation.

OpenStudy (saifoo.khan):

i myself dont know how to solve after solving the square roots.. hehe

OpenStudy (anonymous):

thats why, right. haha. but the answer is correct?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

did u get \[x ^{2}+10x +7\]?

OpenStudy (saifoo.khan):

i have no idea how u got that. ;D

OpenStudy (anonymous):

I did. But try looking at the wolfram link

OpenStudy (anonymous):

\[\sqrt{(x-1)^{2}}<2\sqrt{(x+2)^{2}}\]

OpenStudy (anonymous):

haha this is what i did, do correct me if i'm wrong\[\left( x-1 \right)^{2} < 2 \left( x+2 \right)^{2}\]\[x ^{2}-2x +1 < 2x ^{2}+8x+8\]\[x ^{2}+10x+7 >0\]then i'm all lost

OpenStudy (saifoo.khan):

Where'd the square roots go now?

OpenStudy (aravindg):

ya u should not take away the square root!!

OpenStudy (anonymous):

isn't square root and another square root can be cancel off?

OpenStudy (anonymous):

Yep, and it kinda gives a clue of where the -5 came from but yeah other than that I'm not really sure.

OpenStudy (anonymous):

ok wait, so the -5 came from where again?

OpenStudy (anonymous):

so how do we know how to put that square root and when not to?

OpenStudy (anonymous):

since you can cancel off the square roots and the squares its like a normal equation of \[x-1<2(x+2)\] Not sure Kiki :( kinda confused with some of the rules taught by Mr. Nathan

OpenStudy (anonymous):

i dont really understand what the wolfram site says hahaha @lightchaste

OpenStudy (anonymous):

okay got it! thanks everyone :)

OpenStudy (saifoo.khan):

Wait. how?

OpenStudy (anonymous):

i think there's something to do with the square root thing. \[x-1<2(x+2)\]\[x-1<2x+4\] and \[x-1<-4-2x\]But still the answer is not that accurate. right?

OpenStudy (anonymous):

\[x-1<2x + 4\] \[-x<5\] \[x>-5\] But this isn't accurate either :(

sam (.sam.):

@fatinatikah You're wrong at first, \[(x−1)^2<(2(x+2))^2\] \[x^2-2x+1<4(x^2+4x+4)\] \[x^2-2x+1<4x^2+16x+16\] \[3x^2+18x+15>0\] 3(x^2+6x+5)>0 3(x+1)(x+5)>0 x=-1 , x=-5 Ans x<-5 , x>-1

OpenStudy (anonymous):

:D Awesome! so you have to include the number 2?

sam (.sam.):

yes, you square both sides

OpenStudy (anonymous):

so it isn't just basically converting the absolutes? And how do we determine whether the absolute needs to be converted into squares or with a square root?

sam (.sam.):

just square them

OpenStudy (anonymous):

owh okay :)